A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.00 s for the ball to reach its maximum height. Find (a) the ball’s initial velocity and (b) the height it reaches.
1
Expert's answer
2020-10-09T07:06:48-0400

a) The speed with an equally accelerated movement and the presence of an initial speed is determined by the formula: "V_1=V_0-g\times t" , since the speed at the maximum height is zero, that is, "V_1=0" , then

"0=V_0-g\times t" ;

"V_0=g\times t=10\times 3=30" m/s.

b) Since the distance travelled by the body with equal motion is determined by the formula "S = V _ 0\times t-\frac {g\times t ^ 2} {2}" , the maximum height will be reached at the upper point, therefore

"h=30\times 3-\frac{10\times 3^2}{2}=45" m.